# Learning to Track Any Points from Human Motion

### v1

@@ Abstract
1 Introduction
Accurately tracking points on an object’s surface within a video, a task known as point tracking [14, 22, 28, 12, 15, 27, 16, 53], has become important for various applications, including robotics [4, 60, 69, 66], visual odometry [9, 50], 3D [62, 61] or 4D reconstruction [64, 70, 52, 2, 30, 18], video editing [19, 25], and motion segmentation [23, 29].
Despite its growing significance, acquiring the extensive and diverse data remains a significant challenge. Specifically, manually annotating point trajectories in video sequences is extremely labor-intensive and time-consuming [14, 71], making it a major bottleneck for acquiring real training data. Consequently, the limited availability of real-world training data often restrict the ability of current tracking models to generalize effectively across the diverse real-world scenarios [27, 16]. Although synthetic datasets [41, 21, 71] can be generated at scale, they often fail to capture the complex visual characteristics inherent to real-world scenarios.
Several recent works attempt to address this issue by utilizing self-distillation on real videos [16, 27]; however, these methods often demand extensive computational resources and vast datasets (e.g., 15M videos with 256 GPUs in Doersch et al. [16]) and often show limited improvement, possibly due to the weak supervision signal derived from the model itself. In addition, self-distillation frameworks may suffer from confirmation bias [55]. As a result, efficiently training point tracking models with realistic video data remains an open challenge.
Figure 1: Visualization of videos annotated with our pipeline. In this example, we visualize trajectories extracted using our pipeline using videos from [57]. These examples highlight the complexity of human crowd movement, including both intra-person and inter-person interactions. We leverage this complexity as a source of supervision for training point tracking models.
Our key insight is to leverage the inherent complexities of human motion in real videos by using human model fitting to establish point correspondences across frames, thereby automatically generating complex pseudo-labeled training data for point tracking. For instance, as an individual walks, dances, or turns, this point can be obscured by limbs, or hidden by other articulated bodies in a crowded scene. When multiple people are present, the complexity increases further, as exemplified in Figure 1. In addition, videos capturing human activity reflect this richness, featuring motion blur, varying lighting conditions, and reflective surfaces, all of which are highly difficult to replicate authentically in simulated environments. We believe that these challenges from human motion are essential for training robust point trackers. Furthermore, as shown in [67, 13], point motion carries an inherent 3D structure. Given that 2D trajectories are projections of such 3D motion, leveraging the underlying human 3D geometry provides a natural and effective way to improve 2D tracking accuracy.
To this end, we introduce AnthroTAP, a human-centric pseudo-labeling pipeline for point tracking. AnthroTAP initially fits the SMPL model to humans detected in each video frame using a pre-trained Human Mesh Recovery (HMR) method [17], producing a 3D mesh for each person. The trajectories of these 3D mesh vertices are projected onto the 2D image plane, forming initial pseudo-trajectories. To manage occlusions, we apply a ray-casting leveraging the 3D mesh structure to estimate point visibility, considering both self-occlusion and inter-person occlusion. Finally, to enhance pseudo-label reliability, we introduce a filtering stage based on optical flow consistency [42] between adjacent frames. This stage identifies and removes inaccurate trajectory segments resulting from mesh fitting errors, unmodeled occlusions, or complex non-rigid motions such as loose clothing.
We validate our approach by training a point tracking model using the generated pseudo-labels. The effectiveness of our data generation pipeline is demonstrated by the model’s performance in complex real-world scenarios, achieving state-of-the-art results on the TAP-Vid benchmark [14]. Our method outperforms CoTracker3 [27], which utilizes 11$\times$ more training data, and BootsTAPIR [16], which uses approximately 10,000$\times$ more videos. This underscores the advantage of leveraging complex human motion for efficient and effective point tracking. Moreover, unlike previous works such as BootsTAPIR [16] and CoTracker3 [28] that uses proprietary videos, our dataset is non-proprietary and can be fully open-sourced, contributing valuable resources to the research community.
@@ 1 Introduction
  • Improves the reliability of pseudo-trajectory quality with a filtering stage based on optical flow consistency by identifying and removing inaccuracies caused by mesh fitting errors, unmodeled occlusions, or complex non-rigid motions.
@@ 2 Related Work
PIPs [22] iteratively refines point estimations by constructing local correlation maps, inspired by optical flow methods such as RAFT [58]. TAPIR [15] extends this idea by combining per-frame initialization from TAP-Net [14] with iterative refinement inspired by PIPs. CoTracker [28] learns to track multiple points jointly to enhance reliability. LocoTrack [12] introduces a local 4D correlation to establish better correspondence within videos.
@@
Despite recent advances in model architectures for track refinement, existing methods [22, 14, 15, 12, 37, 36, 48, 11, 34, 1] predominantly rely on synthetic data [41, 21, 71] during training, which can limit the robustness of point tracking in complex real-world scenarios. To alleviate this issue, Chrono [32] leverages representations learned by DINOv2 [45], a model pre-trained on large-scale real-world datasets. Similarly, BootsTAP [16] and CoTracker3 [27] incorporate unlabeled real-world data via self-training to further reduce dependency on synthetic data. However, these methods do not explicitly model the distinctive complexities found in human motion, such as significant non-rigid deformations, diverse textures from clothes, and frequent occlusions arising from interactions within and between individuals. In contrast, our approach specifically targets learning from human motion, which includes these challenging yet commonly occurring real-world phenomena.
@@
Point tracking datasets and benchmarks. Kubric [21] is a data generator that creates synthetic scenes for various tasks [51, 54, 49], including point tracking. While most point tracking methods [15, 28, 37, 12, 34] are trained solely on the Kubric dataset, it features limited motion, primarily consisting of objects falling onto the floor. TAP-Vid [14] provides manual annotations on unlabeled real-world datasets [31, 35, 47]. However, manually annotating point tracks in real videos is extremely time-consuming, making it impractical to scale such efforts for generating training data. DriveTrack [3] explores automated dataset generation in driving scenes [56]; however, the resulting trajectories lack complexity due to relatively simple vehicle motions, no deformation or articulated motion.
@@
Human Mesh Recovery. Goal of Human Mesh Recovery (HMR) is to estimate a full 3D human mesh from a monocular 2D image or video. SMPL [39] proposes an optimization method for parameterized human pose and shape. Early works [6, 33] leverage 2D joint prediction obtained from off-the-shelf 2D pose estimation models [7, 68] and predict 3D body model defined by SMPL. With the advancement of 3D body model prediction from single images, several video-based approaches [17, 20] have been developed. TokenHMR [17] introduces tokenized representation for human pose and achieves state-of-the-art accuracy on multiple in-the-wild 3D benchmarks. We aim to utilize the 3D prior from the Human Mesh Recovery model, to automatically generate complex pseudo-labeled datasets and improve point trackers with the annotated datasets.
@@ 3 Method
The development of robust point tracking models often limited by the scarcity of large-scale, diverse real-world training data. This scarcity stems from the prohibitively time-consuming and labor-intensive nature of manually annotating point trajectories in real-world videos [14], making large-scale annotation infeasible. Although synthetic datasets [41, 21, 71, 26] can be generated at scale, they often fail to capture key characteristics of real-world scenes, such as complex motions and lighting, realistic appearance variations, and dynamic occlusions. Recent efforts have attempted to overcome this limitation using self-training approaches on unlabeled real videos [27, 16]. While promising, these methods suffer from weak supervision signals and often rely on random or heuristic trajectory sampling, limiting their training efficiency.
To address this challenge, we propose AnthroTAP, a pseudo-labeling pipeline that enables leveraging human motion as a natural source of rich supervision for point tracking. Human activities inherently involve complex visual and geometric phenomena, such as non-rigid deformations, articulated motion, occlusions, and diverse clothing appearances, as shown in Figure 1. To benefit from these, we leverage point trajectory across time, using the Skinned Multi-Person Linear (SMPL) model [38], a parametric 3D representation of the human body to generate complex pseudo-labels. Since each vertex on the SMPL mesh corresponds to a fixed anatomical location, it enables consistent and temporally stable point trajectories to be generated from human-centric videos.
To further improve the reliability of the resulting pseudo-labels, we introduce a refinement step based on short-range optical flow consistency, a method proven effective over short temporal windows [11]. This process filters out erroneous trajectory segments caused by mesh fitting errors or occlusions from objects not modeled by SMPL, such as furniture or scene clutter.
@@ 3.1 Point Tracking with Human Mesh Models
Our pseudo-label generation process is founded upon utilizing the SMPL model [38] as a representation for human bodies. For each detected person, indexed by $p$, in a video frame $t$, the SMPL model defines a 3D body mesh $\mathcal{M}_{p,t}$ composed of $N_{v}$ vertices. This mesh is parameterized by low-dimensional 3D human shape parameters $\bm{\beta}_{p}\in\mathbb{R}^{D_{\beta}}$ (typically assumed constant for a person across a sequence) and 3D human pose parameters $\bm{\theta}_{p,t}\in\mathbb{R}^{D_{\theta}}$. $D_{\beta}$ represents the dimensionality of the identity-specific shape space, while $D_{\theta}$ corresponds to the dimensionality of the pose space, capturing body articulation and global orientation. The function $M(\bm{\beta}_{p},\bm{\theta}_{p,t})$ maps these parameters to a set of 3D vertex locations $\{\mathbf{v}_{p,t,j}\in\mathbb{R}^{3}\}_{j=1}^{N_{v}}$.
Figure 2: Overall pipeline. We extract human meshes using an off-the-shelf human mesh recovery model [17], and track points by projecting the mesh vertices. We also determine point visibility using ray-casting. In parallel, we extract optical flow [65] and retain only reliable flow using forward-backward consistency [42]. Finally, we filter the trajectories by checking the consistency between the optical flow and the trajectories generated from the human mesh.
To obtain these SMPL parameters from video frames, we employ a pre-trained video Human Mesh Recovery (HMR) method [17]. Leveraging a strong human prior, the model can reliably track and reconstruct human bodies even under motion blur, extreme motion, or occlusions, which are scenarios where conventional point tracking methods often fail. Given an input video $\{I_{t}\}_{t=1}^{T}$, where $T$ is the number of frames, the HMR model processes the video and outputs the estimated shape parameters $\bm{\beta}_{p}$ and pose parameters $\bm{\theta}_{p,t}$ for each detected person $p$. When the HMR process yields reliable estimations across a sequence of frames, the temporal sequence of corresponding 3D vertex locations $\{\mathbf{v}_{p,t,j}\}_{t\in\mathcal{T}_{p,j}}$ for a specific vertex index $j$ on person $p$ provides an initial 3D trajectory estimation for a point on the human body surface. Here, $\mathcal{T}_{p,j}$ represents the set of frame indices for which vertex $j$ of person $p$ is successfully tracked.
@@
We determine visibility using a ray casting approach within the estimated 3D scene. For each vertex $\mathbf{v}_{p,t,j}$ of person $p$ at frame $t$, we cast a ray from the camera’s optical center $\mathbf{c}$ towards $\mathbf{v}_{p,t,j}$. This ray can be parameterized as $\mathbf{r}(s)=\mathbf{c}+s(\mathbf{v}_{p,t,j}-\mathbf{c})$ for $s\geq 0$. We then identify all intersection points between this ray and the triangular faces of all human meshes $\mathcal{M}_{p^{\prime},t}$ present in frame $t$, considering all persons $p^{\prime}$ from $1$ to $N_{P,t}$. Let $s_{\text{int},k}$ be the parameter $s$ for the $k$-th intersection point. The vertex $\mathbf{v}_{p,t,j}$ is considered visible ($v_{p,t,j}=1$) if and only if the minimum $s$ value among all such intersection points is greater than or equal to $1$. This implies that no part of any modeled human mesh intersects the ray segment $(\mathbf{c},\mathbf{v}_{p,t,j})$ strictly before $\mathbf{v}_{p,t,j}$ itself. Formally:
$v_{p,t,j}=\mathbbm{1}\left[\min_{k,\forall p^{\prime}}\{s_{\text{int},k}\mid\text{intersection occurs with any mesh }\mathcal{M}_{p^{\prime},t}\}\geq 1\right].$ (1)

### v2

# AnthroTAP: Learning Point Tracking with Real-World Motion

@@ Abstract
Point tracking models often struggle to generalize to real-world videos because large-scale training data is predominantly synthetic–the only source currently feasible to produce at scale. Collecting real-world annotations, however, is prohibitively expensive, as it requires tracking hundreds of points across frames. We introduce AnthroTAP, an automated pipeline that generates large-scale pseudo-labeled point tracking data from real human motion videos. Leveraging the structured complexity of human movement-non-rigid deformations, articulated motion, and frequent occlusions–AnthroTAP fits Skinned Multi-Person Linear (SMPL) models to detected humans, projects mesh vertices onto image planes, resolves occlusions via ray-casting, and filters unreliable tracks using optical flow consistency. A model trained on the AnthroTAP dataset achieves state-of-the-art performance on TAP-Vid, a challenging general-domain benchmark for tracking any point on diverse rigid and non-rigid objects (e.g., humans, animals, robots, and vehicles). Our approach outperforms recent self-supervised teacher-student models trained on vastly larger real datasets, while requiring only one day of training on 4 GPUs. AnthroTAP shows that structured human motion offers a scalable and effective source of real-world supervision for point tracking. Code and datasets will be made publicly available.
1 Introduction
Figure 1: Visualization of videos annotated with our pipeline. In this example, we visualize trajectories extracted using our pipeline using videos from [peize2021dance]. These examples highlight the complexity of human crowd movement, including both intra-person and inter-person interactions. We leverage this complexity as a source of supervision for training point tracking models.
Accurately tracking points on an object’s surface within a video, a task known as point tracking [doersch2022tap, harley2022particle, karaev2024cotracker, cho2024local, doersch2023tapir, karaev2024cotracker3, doersch2024bootstap, shrivastava2024gmrw], has become important for various applications, including robotics [bharadhwaj2024track2act, vecerik2024robotap, yang2025magma, wen2024anypointtrajectorymodelingpolicy], visual odometry [chen2024leap, rockwell2025dynpose], 3D [wang2025vggt, wang2024vggsfm] or 4D reconstruction [wang2024shape, seidenschwarz2024dynomo, badki2026l4p, kastenfast, st4rtrack2025], video editing [geng2024motion, jeong2024track4gen], and motion segmentation [huang2025segmentmotionvideos, karazija24learning].
Despite its growing significance, acquiring the extensive and diverse real-world data remains a significant challenge. Specifically, manually annotating point trajectories in video sequences is extremely labor-intensive and time-consuming [doersch2022tap, zheng2023pointodyssey], making it a major bottleneck for acquiring real training data. Consequently, the limited availability of real-world training data often restrict the ability of current tracking models to generalize effectively across the diverse real-world scenarios [karaev2024cotracker3, doersch2024bootstap]. Although synthetic datasets [mayer2016large, greff2022kubric, zheng2023pointodyssey] can be generated at scale, they often fail to capture the complex visual characteristics inherent to real-world scenarios.
Several recent works attempt to address this issue by utilizing self-distillation on real videos [doersch2024bootstap, karaev2024cotracker3]; however, these methods often demand extensive computational resources and vast datasets (e.g., 15M videos with 256 GPUs in Doersch et al. [doersch2024bootstap]) and often show limited improvement, possibly due to the weak supervision signal derived from the model itself. In addition, self-distillation frameworks may suffer from confirmation bias [sohn2020fixmatch]. As a result, efficiently training point tracking models with realistic video data remains an open challenge.
Our key insight is that human motion offers a unique combination of properties: by distilling knowledge from human mesh fitting models, we can automatically establish reliable point correspondences, leveraging real-world data characteristics and inherent complexities such as non-rigid deformations, articulated movements, and frequent occlusions to generate complex and effective pseudo-labeled training data for point tracking without manual annotation. For instance, as an individual walks, dances, or turns, this point can be obscured by limbs, or hidden by other articulated bodies in a crowded scene. When multiple people are present, the complexity increases further, as exemplified in Figure 1. In addition, videos capturing human activity reflect this real-world richness, featuring motion blur, varying lighting conditions, and reflective surfaces, all of which are highly difficult to replicate authentically in simulated environments. We believe that these challenges from human motion with real-world visual characteristics provides essential training signals for robust trackers. Furthermore, as shown in [Xiao_2024_CVPR, cho2025seurat], point motion carries an inherent 3D structure. Given that 2D trajectories are projections of such 3D motion, leveraging the underlying human 3D geometry provides a natural and effective way to improve 2D tracking accuracy.
To this end, we introduce AnthroTAP, a pseudo-labeling pipeline that distills knowledge from human motion to automatically generate real-world point tracking training data. AnthroTAP initially fits the SMPL model to humans detected in each video frame using a pre-trained Human Mesh Recovery (HMR) method [dwivedi2024tokenhmr], producing a 3D mesh for each person. The trajectories of these 3D mesh vertices are projected onto the 2D image plane, forming initial pseudo-trajectories. To manage occlusions, we apply a ray-casting leveraging the 3D mesh structure to estimate point visibility, considering both self-occlusion and inter-person occlusion. Finally, to enhance pseudo-label reliability, we introduce a filtering stage based on optical flow consistency [meister2018unflow] between adjacent frames.
@@ 1 Introduction
  • Introduces a simple yet effective approach to leverage real-world training data for point tracking by distilling knowledge from human motion, enabling automatic generation of pseudo-labeled datasets with strong supervision signals that capture complex motion patterns and real-world visual characteristics.
@@
Figure 2: Overall pipeline. We extract human meshes using an off-the-shelf human mesh recovery model [dwivedi2024tokenhmr], and track points by projecting the mesh vertices. We also determine point visibility using ray-casting. In parallel, we extract optical flow [wang2024sea] and retain only reliable flow using forward-backward consistency [meister2018unflow]. Finally, we filter the trajectories by checking the consistency between the optical flow and the trajectories generated from the human mesh.
@@ 2 Related work
PIPs [harley2022particle] iteratively refines point estimations by constructing local correlation maps, inspired by optical flow methods such as RAFT [teed2020raft]. TAPIR [doersch2023tapir] extends this idea by combining per-frame initialization from TAP-Net [doersch2022tap] with iterative refinement inspired by PIPs. CoTracker [karaev2024cotracker] learns to track multiple points jointly to enhance reliability. LocoTrack [cho2024local] introduces a local 4D correlation to establish better correspondence within videos.
@@
Despite recent advances in model architectures for track refinement, existing methods [harley2022particle, doersch2022tap, doersch2023tapir, cho2024local, li2024taptr, li2024taptrv2, qu2024taptrv3, cho2024flowtrack, lemoing2024dense, aydemir2025track] predominantly rely on synthetic data [mayer2016large, greff2022kubric, zheng2023pointodyssey] during training, which can limit the robustness of point tracking in complex real-world scenarios. To alleviate this issue, Chrono [kim2025exploring] leverages representations learned by DINOv2 [oquab2023dinov2], a model pre-trained on large-scale real-world datasets. Similarly, BootsTAP [doersch2024bootstap] and CoTracker3 [karaev2024cotracker3] incorporate unlabeled real-world data via self-training to further reduce dependency on synthetic data. However, these methods do not explicitly model the distinctive complexities found in human motion, such as significant non-rigid deformations, diverse textures from clothes, and frequent occlusions arising from interactions within and between individuals. In contrast, our approach specifically targets learning from human motion, which includes these challenging yet commonly occurring real-world phenomena.
@@
Point tracking datasets and benchmarks. Kubric [greff2022kubric] is a data generator that creates synthetic scenes for various tasks [sajjadi2022scene, singh2022simple, ren2025gen3c], including point tracking. While most point tracking methods [doersch2023tapir, karaev2024cotracker, li2024taptr, cho2024local, lemoing2024dense] are trained solely on the Kubric dataset, it features limited motion, primarily consisting of objects falling onto the floor. TAP-Vid [doersch2022tap] provides manual annotations on unlabeled real-world datasets [kay2017kinetics, lee2021beyond, pont20172017]. However, manually annotating point tracks in real videos is extremely time-consuming, making it impractical to scale such efforts for generating training data. DriveTrack [balasingam2024drivetrack] explores automated dataset generation in driving scenes [Sun_2020_CVPR]; however, the resulting trajectories lack complexity due to relatively simple vehicle motions, no deformation or articulated motion.
@@
Human Mesh Recovery. Goal of Human Mesh Recovery (HMR) is to estimate a full 3D human mesh from a monocular 2D image or video. SMPL [loper2023smpl] proposes an optimization method for parameterized human pose and shape. Early works [bogo2016keep, lassner2017unite] leverage 2D joint prediction obtained from off-the-shelf 2D pose estimation models [cao2017realtime, xu2022vitpose] and predict 3D body model defined by SMPL. With the advancement of 3D body model prediction from single images, several video-based approaches [dwivedi2024tokenhmr, goel2023humans] have been developed. TokenHMR [dwivedi2024tokenhmr] introduces tokenized representation for human pose and achieves state-of-the-art accuracy on multiple in-the-wild 3D benchmarks. We aim to utilize the 3D prior from the Human Mesh Recovery model, to automatically generate complex pseudo-labeled datasets and improve point trackers with the annotated datasets.
@@ 3 Method
The development of robust point tracking models often limited by the scarcity of large-scale, diverse real-world training data. This scarcity stems from the prohibitively time-consuming and labor-intensive nature of manually annotating point trajectories in real-world videos [doersch2022tap], making large-scale annotation infeasible. Although synthetic datasets [mayer2016large, greff2022kubric, zheng2023pointodyssey, karaev2023dynamicstereo] can be generated at scale, they often fail to capture key characteristics of real-world scenes, such as complex motions and lighting, realistic appearance variations, and dynamic occlusions. Recent efforts have attempted to overcome this limitation using self-training approaches on unlabeled real videos [karaev2024cotracker3, doersch2024bootstap]. While promising, these methods suffer from weak supervision signals and often rely on random or heuristic trajectory sampling, limiting their training efficiency.
To address this challenge, we propose AnthroTAP, a pseudo-labeling pipeline that enables leveraging human motion as a natural source of rich supervision for point tracking. Human activities inherently involve complex visual and geometric phenomena, such as non-rigid deformations, articulated motion, occlusions, and diverse clothing appearances, as shown in Figure 1. To benefit from these, we leverage point trajectory across time, using the Skinned Multi-Person Linear (SMPL) model [SMPL:2015], a parametric 3D representation of the human body to generate complex pseudo-labels. Since each vertex on the SMPL mesh corresponds to a fixed anatomical location, it enables consistent and temporally stable point trajectories to be generated from human-centric videos.
To further improve the reliability of the resulting pseudo-labels, we introduce a refinement step based on short-range optical flow consistency, a method proven effective over short temporal windows [cho2024flowtrack]. This process filters out erroneous trajectory segments caused by mesh fitting errors or occlusions from objects not modeled by SMPL, such as furniture or scene clutter.
@@ 3.1 Point tracking with human mesh models
Our pseudo-label generation process is founded upon utilizing the SMPL model [SMPL:2015] as a representation for human bodies. For each detected person, indexed by $p$, in a video frame $t$, the SMPL model defines a 3D body mesh $\mathcal{M}_{p,t}$ composed of $N_{v}$ vertices. This mesh is parameterized by low-dimensional 3D human shape parameters $\bm{\beta}_{p}\in\mathbb{R}^{D_{\beta}}$ (typically assumed constant for a person across a sequence) and 3D human pose parameters $\bm{\theta}_{p,t}\in\mathbb{R}^{D_{\theta}}$. $D_{\beta}$ represents the dimensionality of the identity-specific shape space, while $D_{\theta}$ corresponds to the dimensionality of the pose space, capturing body articulation and global orientation. The function $M(\bm{\beta}_{p},\bm{\theta}_{p,t})$ maps these parameters to a set of 3D vertex locations $\{\mathbf{v}_{p,t,j}\in\mathbb{R}^{3}\}_{j=1}^{N_{v}}$.
To obtain these SMPL parameters from video frames, we employ a pre-trained video Human Mesh Recovery (HMR) method [dwivedi2024tokenhmr]. Leveraging a strong human prior, the model can reliably track and reconstruct human bodies even under motion blur, extreme motion, or occlusions, which are scenarios where conventional point tracking methods often fail. Given an input video $\{I_{t}\}_{t=1}^{T}$, where $T$ is the number of frames, the HMR model processes the video and outputs the estimated shape parameters $\bm{\beta}_{p}$ and pose parameters $\bm{\theta}_{p,t}$ for each detected person $p$. When the HMR process yields reliable estimations across a sequence of frames, the temporal sequence of corresponding 3D vertex locations $\{\mathbf{v}_{p,t,j}\}_{t\in\mathcal{T}_{p,j}}$ for a specific vertex index $j$ on person $p$ provides an initial 3D trajectory estimation for a point on the human body surface. Here, $\mathcal{T}_{p,j}$ represents the set of frame indices for which vertex $j$ of person $p$ is successfully tracked.
@@
We determine visibility using a ray casting approach within the estimated 3D scene. For each vertex $\mathbf{v}_{p,t,j}$ of person $p$ at frame $t$, we cast a ray from the camera’s optical center $\mathbf{c}$ towards $\mathbf{v}_{p,t,j}$. This ray can be parameterized as $\mathbf{r}(s)=\mathbf{c}+s(\mathbf{v}_{p,t,j}-\mathbf{c})$ for $s\geq 0$. We then identify all intersection points between this ray and the triangular faces of all human meshes $\mathcal{M}_{p^{\prime},t}$ present in frame $t$, considering all persons $p^{\prime}$ from $1$ to $N_{P,t}$. Let $s_{\text{int},k}$ be the parameter $s$ for the $k$-th intersection point. The vertex $\mathbf{v}_{p,t,j}$ is considered visible ($v_{p,t,j}=1$) if and only if the minimum $s$ value among all such intersection points is greater than or equal to $1$. This implies that no part of any modeled human mesh intersects the ray segment $(\mathbf{c},\mathbf{v}_{p,t,j})$ strictly before $\mathbf{v}_{p,t,j}$ itself.